# Limited Memory LRSGA Optimizer to competitive optimization

### v1

@@ 3 Limited Memory LRSGA
3:  $M_{k+1}\leftarrow M_{k}+\frac{\bigl(y_{k}^{f}-M_{k}\,s_{k}^{y}\bigr)}{(s_{k}^{w})^{\top}s_{k}^{w}}\,(s_{k}^{y})^{\top}$
4:  $N_{k+1}\leftarrow N_{k}+\frac{\bigl(y_{k}^{g}-N_{k}\,s_{k}^{x}\bigr)}{(s_{k}^{w})^{\top}s_{k}^{w}}\,(s_{k}^{x})^{\top}$
5:  $x_{k+1}\leftarrow x_{k}-\eta\Bigl(\partial_{x}f(w_{k})-\frac{\tau}{2}\bigl(M_{k}\,-N_{k}^{\top}\,\bigr)\Bigr)\partial_{y}g_{k}$
6:  $y_{k+1}\leftarrow y_{k}-\eta\Bigl(\partial_{y}g(w_{k})-\frac{\tau}{2}\bigl(N_{k}\,-M_{k}^{\top}\,\bigr)\partial_{x}f_{k}\Bigr)$
@@ 3.1 Limited-memory recursive updates for $M_{k}$ and $N_{k}$
Consequently, iterations 8 can be rewrite as:
@@
$M_{k}$ $=M_{k-\ell}(\widetilde{V}_{k-\ell}\dots\widetilde{V}_{k-1})+p_{k-\ell}y_{k-\ell}^{f}(s_{k-\ell}^{y})^{\top}(\widetilde{V}_{k-\ell+1}\dots\widetilde{V}_{k-1})+\dots$
@@
$N_{k}$ $=N_{k-\ell}({V}_{k-\ell}\dots V_{k-1})+p_{k-\ell}y_{k-\ell}^{g}(s_{k-\ell}^{x})^{\top}({V}_{k-\ell+1}\dots V_{k-1})+\dots$
@@
$M_{k}^{\top}=\Bigl(\widetilde{V}_{k-1}\cdots\widetilde{V}_{k-\ell}\Bigr)(M_{(k-\ell)})^{\top}+p_{k-\ell}^{y}\Bigl(\widetilde{V}_{k-1}\cdots\widetilde{V}_{k-\ell+1}\Bigr)s_{k-\ell}^{y}\,(y_{k-\ell}^{f})^{\top}+\cdots+p_{k-1}^{y}\,s_{k-1}^{y}\,(y_{k-1}^{f})^{\top},$
@@
$N_{k}^{\top}=\Bigl(V_{k-1}\cdots V_{k-\ell}\Bigr)(N_{(k-\ell)})^{\top}+p_{k-\ell}^{x}\Bigl(V_{k-1}\cdots V_{k-\ell+1}\Bigr)s_{k-\ell}^{x}\,(y_{k-\ell}^{g})^{\top}+\cdots+p_{k-1}^{x}\,s_{k-1}^{x}\,(y_{k-1}^{g})^{\top}.$
@@ 3.2 Adapted two-loop recursion algorithms
Classical L-BFGS uses the two–loop recursion to apply a square (inverse) Hessian approximation $H_{k}\in\mathbb{R}^{d\times d}$ to a vector. In contrast, our setting requires multiplying rectangular mixed-Jacobian approximation and their transpositions, $M_{k}\in\mathbb{R}^{m\times n}$ and $N_{k}\in\mathbb{R}^{n\times m}$, by vectors of compatible size. To this end, we adapt the recursion by storing the $\ell$ pairs $\{(s_{t}^{y},y_{t}^{f})\}$ and $\{(s_{t}^{x},y_{t}^{g})\}$, and by using side-specific projectors $\widetilde{V}_{t}:=I-p_{t}s_{t}^{y}(s_{t}^{y})^{\top}$ and $V_{t}:=I-p_{t}s_{t}^{x}(s_{t}^{x})^{\top}$, with the common scaling $p_{t}:=1/\bigl\|(s_{t}^{w})\bigr\|_{2}^{2}$. This preserves the appropriate secant conditions for the mixed blocks while avoiding storage of the full matrices.
@@ 3.2.1 Direct Two-loop recursion
1:Input: $\{p_{k-1},p_{k-2},\dots,p_{k-\ell}\},\{s_{k-1}^{j},s_{k-2}^{j},\dots,s_{k-\ell}^{j}\},\{y_{k-1}^{h},y_{k-2}^{h},\dots,y_{k-\ell}^{h}\}$, $j\in\{x,y\}$, $h\in\{f,g\}$, vettore $q$, matrice iniziale $H_{0}^{k}$
@@
$\boxed{\;M_{k}\,\partial_{y}g(w_{k})\;=\;\textbf{TwoLoopRecursionDirect-K}\bigl(\{p_{t},s_{t}^{y},y_{t}^{f}\}_{t=k-\ell}^{k-1},\ q=\partial_{y}g(w_{k}),\ H_{0}^{k}=M_{k-\ell}\bigr),\;}$
$\boxed{\;N_{k}\,\partial_{x}f(w_{k})\;=\;\textbf{TwoLoopRecursionDirect-K}\bigl(\{p_{t},s_{t}^{x},y_{t}^{g}\}_{t=k-\ell}^{k-1},\ q=\partial_{x}f(w_{k}),\ H_{0}^{k}=N_{k-\ell}\bigr).\;}$
@@ 3.2.2 Transpose Two-loop recursion
1:Input: $\{p_{k-1},p_{k-2},\dots,p_{k-\ell}\},\{s_{k-1}^{j},s_{k-2}^{j},\dots,s_{k-\ell}^{j}\},\{y_{k-1}^{h},y_{k-2}^{h},\dots,y_{k-\ell}^{h}\}$, $j\in\{x,y\}$, $h\in\{f,g\}$, vector $q$, initial matrix $(H_{0}^{k})^{\top}$
@@
$\boxed{\;M_{k}^{\top}\,\partial_{x}f(w_{k})\;=\;\textbf{TwoLoopRecursionTranspose-K}\bigl(\{p_{t},s_{t}^{y},y_{t}^{f}\}_{t=k-\ell}^{k-1},\ q=\partial_{x}f(w_{k}),\ H_{0}^{k}=(M_{k-\ell})^{\top}\ \bigr),\;}$
$\boxed{\;N_{k}^{\top}\,\partial_{y}g(w_{k})\;=\;\textbf{TwoLoopRecursionTranspose-K}\bigl(\{p_{t},s_{t}^{x},y_{t}^{g}\}_{t=k-\ell}^{k-1},\ q=\partial_{y}g(w_{k}),\ H_{0}^{k}=(N_{k-\ell})^{\top}\bigr).\;}$
Finally, to avoid storing a full base matrix $H_{k}^{0}$ (and its transpose) in the two–loop recursions, we set
$H_{0}^{k}:=\frac{y_{k-\ell-1}^{\,h}\,(s_{k-\ell-1}^{\,j})^{\top}}{(s_{k-\ell-1})^{\top}(s_{k-\ell-1})},$
where $j\in\{x,y\}$ an $h\in\{f,g\}$ matches the recursion (i.e., $j=y$ and $h=f$ for the $M_{k}$-recursion, and $j=x$, $h=g$ for the $N_{k}$-recursion and $s_{t}=(s^{x}_{t},s^{y}_{t})$ is the joint vector. With this rank–one choice, matrix–vector products reduce to inner products:
$H_{k}^{0}q\;=\;\frac{y_{k-1}^{\,h}\,\big((s_{k-1}^{\,j})^{\top}q\big)}{(s_{k-\ell-1})^{\top}(s_{k-\ell-1})},\qquad(H_{k}^{0})^{\top}q\;=\;\frac{s_{k-1}^{\,j}\,\big((y_{k-1}^{\,h})^{\top}q\big)}{(y_{k-1}^{h})^{\top}(y_{k-1}^{h})}.$
We remark that in both differences $\partial_{x}f(w_{k+1})-\partial_{x}f(w_{k})$ and $\partial_{y}g(w_{k+1})-\partial_{y}g(w_{k})$ relate two consecutive iterates and this implies that each term is evaluated on the current mini–batch $k$ (e.g., $\partial_{x}f_{B_{k+1}}(w_{k+1})-\partial_{x}f_{B_{k}}(w_{k})$). We emphasize that this fact degrades the accuracy of the mixed–second–derivative approximations used by our secant updates and so a stochastic implementation is required.
@@
We propose an additional variant, LM–LRSGA–EMA, which mitigates mini-batch–induced noise without manipulating the batches. Specifically, we maintain exponential moving averages of $y_{k}^{\,f}$ and $y_{k}^{\,g}$ and use them in the two–loop recursions:

### v2

@@ 4 Limited Memory LRSGA
3:  $M_{k}\leftarrow M_{k-1}+\frac{\bigl(y_{k-1}^{f}-M_{k-1}\,s_{k-1}^{y}\bigr)}{(s_{k-1}^{w})^{\top}s_{k-1}^{w}}\,(s_{k-1}^{y})^{\top}$
4:  $N_{k}\leftarrow N_{k-1}+\frac{\bigl(y_{k-1}^{g}-N_{k-1}\,s_{k-1}^{x}\bigr)}{(s_{k-1}^{w})^{\top}s_{k-1}^{w}}\,(s_{k-1}^{x})^{\top}$
5:  $x_{k+1}\leftarrow x_{k}-\eta\Bigl(\partial_{x}f_{k}-\frac{\tau}{2}\bigl(M_{k}\,-N_{k}^{\top}\,\bigr)\partial_{y}g_{k}\Bigr)$
6:  $y_{k+1}\leftarrow y_{k}-\eta\Bigl(\partial_{y}g_{k}-\frac{\tau}{2}\bigl(N_{k}\,-M_{k}^{\top}\,\bigr)\partial_{x}f_{k}\Bigr)$
@@ 4.1 Limited-memory recursive updates for $M_{k}$ and $N_{k}$
Consequently, iterations 13 can be rewritten as:
@@
$M_{k}^{(\ell)}$ $=M_{k-\ell}(\widetilde{V}_{k-\ell}\dots\widetilde{V}_{k-1})+p_{k-\ell}y_{k-\ell}^{f}(s_{k-\ell}^{y})^{\top}(\widetilde{V}_{k-\ell+1}\dots\widetilde{V}_{k-1})+\dots$ (15)
@@
$N_{k}^{(\ell)}$ $=N_{k-\ell}({V}_{k-\ell}\dots V_{k-1})+p_{k-\ell}y_{k-\ell}^{g}(s_{k-\ell}^{x})^{\top}({V}_{k-\ell+1}\dots V_{k-1})+\dots$ (16)
In practice, we avoid storing the base matrices $M_{k-\ell}$ and $N_{k-\ell}$, by replacing them with respectively the rank-one initializations:
$H_{0}^{k}:=p_{k-\ell}\,y_{k-\ell}^{\,h}\,(s_{k-\ell}^{\,j})^{\top},$ (17)
where $j\in\{x,y\}$ an $h\in\{f,g\}$ matches the recursion (i.e., $j=y$ and $h=f$ for the $M_{k}$-recursion, and $j=x$, $h=g$ for the $N_{k}$-recursion.
Then we can redefine the limited-memory approximations $M_{k}^{(\ell)}$ and $N_{k}^{(\ell)}$ as
$M_{k}^{(\ell)}$ $=H_{0}^{k}(\widetilde{V}_{k-\ell}\dots\widetilde{V}_{k-1})+p_{k-\ell}y_{k-\ell}^{f}(s_{k-\ell}^{y})^{\top}(\widetilde{V}_{k-\ell+1}\dots\widetilde{V}_{k-1})+\dots$ (18)
$N_{k}^{(\ell)}$ $=H_{0}^{k}({V}_{k-\ell}\dots V_{k-1})+p_{k-\ell}y_{k-\ell}^{g}(s_{k-\ell}^{x})^{\top}({V}_{k-\ell+1}\dots V_{k-1})+\dots$ (19)
@@
$(M_{k}^{(\ell)})^{\top}=\Bigl(\widetilde{V}_{k-1}\cdots\widetilde{V}_{k-\ell}\Bigr)(H_{0}^{k})^{\top}+p_{k-\ell}\Bigl(\widetilde{V}_{k-1}\cdots\widetilde{V}_{k-\ell+1}\Bigr)s_{k-\ell}^{y}\,(y_{k-\ell}^{f})^{\top}+\cdots+p_{k-1}\,s_{k-1}^{y}\,(y_{k-1}^{f})^{\top},$
@@
$(N_{k}^{(\ell)})^{\top}=\Bigl(V_{k-1}\cdots V_{k-\ell}\Bigr)(H_{0}^{k})^{\top}+p_{k-\ell}\Bigl(V_{k-1}\cdots V_{k-\ell+1}\Bigr)s_{k-\ell}^{x}\,(y_{k-\ell}^{g})^{\top}+\cdots+p_{k-1}\,s_{k-1}^{x}\,(y_{k-1}^{g})^{\top}.$
@@
Remark 4.2 (Choice of the rank-one base $H_{0}^{k}$).
In the limited-memory implementation we maintain a buffer of length $\ell$ containing the most recent curvature pairs, i.e., indices $t\in\{k-\ell,\dots,k-1\}$ once the buffer is full. Accordingly, in (17) we choose the rank-one base matrix $H_{0}^{k}$ using the oldest pair within the current window, namely $t=k$-$\ell$. With this choice, the oldest curvature pair $(s_{k-\ell}^{j},y_{k-\ell}^{h})$ appears twice in the explicit expansions (18)–(19): once through the base term $H_{0}^{k}(\widetilde{V}_{k-\ell}\cdots\widetilde{V}_{k-1})$ (respectively $H_{0}^{k}(V_{k-\ell}\cdots V_{k-1})$), and once through the first rank-one term in the truncated sum. As an alternative, one may initialize the base using the curvature pair immediately preceding the memory window,
$\hat{H}_{0}^{k}:=p_{k-\ell-1}\,y_{k-\ell-1}^{\,h}\,(s_{k-\ell-1}^{\,j})^{\top}.$
In that case one recovers the telescoping identity $M_{k}-M_{k}^{(\ell)}=M_{k-\ell-1}\,(\widetilde{V}_{k-\ell-1}\cdots\widetilde{V}_{k-1})$, $N_{k}-N_{k}^{(\ell)}=N_{k-\ell-1}\,(V_{k-\ell-1}\cdots V_{k-1}),$ and therefore the sharper bounds $\|M_{k}-M_{k}^{(\ell)}\|_{2}\leq\|M_{k-\ell-1}\|_{2}$ and $\|N_{k}-N_{k}^{(\ell)}\|_{2}\leq\|N_{k-\ell-1}\|_{2}$ (using $\|\widetilde{V}_{t}\|_{2}\leq 1$ and $\|V_{t}\|_{2}\leq 1$). From an implementation standpoint, this alternative requires computing $\hat{H}_{0}^{k}$ before the $(k-\ell-1)$-th curvature pair is discarded, i.e., prior to updating the history lists (i.e., before the pop operation). In this paper (and in our experiments) we adopt the implementation-consistent choice $H_{0}^{k}$ (17) with index $k$-$\ell$, so that no curvature pair beyond the last $\ell$ needs to be accessed or stored. All subsequent norm bounds and local convergence arguments only require a rank-one base of uniformly bounded norm, hence they apply unchanged to either initialization.
@@ 4.2 Adapted two-loop recursion algorithms
Classical L-BFGS uses the two–loop recursion to apply a square (inverse) Hessian approximation $H_{k}\in\mathbb{R}^{d\times d}$ to a vector. In contrast, our setting requires multiplying rectangular mixed-Jacobian approximation and their transpositions, $M_{k}\in\mathbb{R}^{m\times n}$ and $N_{k}\in\mathbb{R}^{n\times m}$, by vectors of compatible size. To this end, we introduced the two following two-loop recursion algorithms.
@@ 4.2.1 Direct Two-loop recursion
1:Input: $\{p_{k-1},p_{k-2},\dots,p_{k-\ell}\},\{s_{k-1}^{j},s_{k-2}^{j},\dots,s_{k-\ell}^{j}\},\{y_{k-1}^{h},y_{k-2}^{h},\dots,y_{k-\ell}^{h}\}$, $j\in\{x,y\}$, $h\in\{f,g\}$, vector $q$, initial matrix $H_{0}^{k}$
@@
$\boxed{\;M_{k}\,\partial_{y}g(w_{k})\;=\;\textbf{TwoLoopRecursionDirect-K}\bigl(\{p_{t},s_{t}^{y},y_{t}^{f}\}_{t=k-\ell}^{k-1},\ q=\partial_{y}g(w_{k}),\ H_{0}^{k}\bigr),\;}$
$\boxed{\;N_{k}\,\partial_{x}f(w_{k})\;=\;\textbf{TwoLoopRecursionDirect-K}\bigl(\{p_{t},s_{t}^{x},y_{t}^{g}\}_{t=k-\ell}^{k-1},\ q=\partial_{x}f(w_{k}),\ H_{0}^{k}\bigr).\;}$
@@ 4.2.2 Transpose Two-loop recursion
1:Input: $\{p_{k-1},p_{k-2},\dots,p_{k-\ell}\},\{s_{k-1}^{j},s_{k-2}^{j},\dots,s_{k-\ell}^{j}\},\{y_{k-1}^{h},y_{k-2}^{h},\dots,y_{k-\ell}^{h}\}$, $j\in\{x,y\}$, $h\in\{f,g\}$, vector $q$, initial matrix $H_{0}^{k}$
@@
$\boxed{\;M_{k}^{\top}\,\partial_{x}f(w_{k})\;=\;\textbf{TwoLoopRecursionTranspose-K}\bigl(\{p_{t},s_{t}^{y},y_{t}^{f}\}_{t=k-\ell}^{k-1},\ q=\partial_{x}f(w_{k}),\ H_{0}^{k}\bigr),\;}$
$\boxed{\;N_{k}^{\top}\,\partial_{y}g(w_{k})\;=\;\textbf{TwoLoopRecursionTranspose-K}\bigl(\{p_{t},s_{t}^{x},y_{t}^{g}\}_{t=k-\ell}^{k-1},\ q=\partial_{y}g(w_{k}),\ H_{0}^{k})^{\top}\bigr).\;}$
We remark that in both differences $\partial_{x}f(w_{k+1})-\partial_{x}f(w_{k})$ and $\partial_{y}g(w_{k+1})-\partial_{y}g(w_{k})$ relate two consecutive iterates and this implies that each term is evaluated on the current mini–batch $B$ (e.g., $\partial_{x}f_{B_{k+1}}(w_{k+1})-\partial_{x}f_{B_{k}}(w_{k})$). We emphasize that this fact degrades the accuracy of the mixed–second–derivative approximations used by our secant updates and so a stochastic implementation is required.
@@
Since common/consistent-batch gradient displacement can increase implementation complexity (and may require additional gradient evaluations at $w_{k+1}$ on the previous mini-batch $B_{k}$), we propose an additional variant, LM-LRSGAEma, which mitigates mini-batch–induced noise without manipulating the batches. Specifically, we maintain exponential moving averages of $y_{k}^{\,f}$ and $y_{k}^{\,g}$ and use them in the two–loop recursions: